f(x)=x^3-x^2+x-1 g(x)=-1/2x

find: (gof)(0)

Answer by stanbon(75887)

f(x)=x^3-x^2+x-1 g(x)=-1/2x
find: (gof)(0) = g[f(0)] = g[0^3-0^2+0-1] = g[-1] = -(1/2)(-1) = 1/2
------------------
Cheers,
Stan H.
-----------